# Gas supply shocks, Uncertainty and Price setting: Evidences from Italian Firms

### v3

@@ 2 October 2022 Russia halts gas deliveries to Italy trough the TAG pipeline
We end up isolating 15 surprises which are reported in Table 1. We construct a daily instrument which takes value of 1 in correspondence of any of the surprises. Then, we aggregate the instrument at the monthly level.
@@
We specify a monthly VAR including six variables. We take the day ahead Dutch TTF as a measure of the spot price of natural gas. Then, in the spirit of adolfsen2024gas, we include the Italian gas quantities, gas inventories and industrial production. We also control for the world industrial production and for the Prezzo Unico Nazionale (PUN) as a measure of the Italian spot electricity price. [^6] Data about gas quantities and gas inventories are provided by the Italian Ministero dell’ambiente e della sicurezza energetica. Gas quantities are defined as national production, plus imports minus exports. Gas inventories are given by the cumulative sum of the variation in inventories, given a base value of 7000 millions of standard cubic meters. Both gas quantities and gas inventories are seasonally adjusted by using the X-13 ARIMA seasonal adjustment method. Data about the PUN are provided by Gestore Mercati Energetici. World industrial production is provided by BH2019. Monitoring the response of the PUN to a gas supply shock is informative because, in a uniform-price electricity market with gas-fired units often setting the marginal cost, like the Italian one, changes in the TTF day-ahead price feed into generators’ bids and hence the Italian spot price with short lags; the PUN therefore reflects downstream pass-through from gas to electricity price. Controlling for world industrial production purges the gas-price equation of global demand pressure, sharpening identification of a gas supply shock. The VAR is estimated in log levels with a constant from 2004M1 until 2025M6. The lag order is set at 12. We assess instrument relevance by performing an F-test in the regression of the TTF residuals from the VAR on the instrument, following olea2021inference. An F-statistic above 10 indicates a strong instrument. In our case, the F-statistic equals 13.6, confirming high instrument strength.
Figure 5: Gas supply shock. The figure shows the natural gas supply shock obtained from the IV-VAR (left) and the corresponding correlogram (right). Sample: 2005:M1–2025:m6.
We also conduct the F-test allowing for heteroskedasticity, which delivers an F-statistic of 15.18.777The instrument explains around 4.9% of the TTF residual. Figure 5 reports the structural shock extracted from the IV-VAR and the corresponding correlogram.
@@
We test the exogeneity of our shock with respect to oil and energy-related shocks identified in the literature using a Wald test of joint significance. Specifically, we regress our shock measure on the contemporaneous and 12 lagged values of each structural innovation considered, as well as on 12 of its own lags. We consider the oil price news shock of Kanzig2020, the carbon price shock of kanzig2023unequal, the oil demand shocks of CALDARA20191 and BH2019, the oil supply,inventories and demand shock of BH2019, the oil price expectation shock of Baumeister2023, and the oil supply shocks of CALDARA20191, kilian2009not and kilian2008exogenous. Results are reported in Table 2. The Wald tests do not reject the null of no contamination for any proxy, supporting their exogeneity. Only the BH2019 oil inventories shock approaches significance at the 10% level.
@@
Kanzig2020 Oil price news 2000Q4 - 2024Q2 7.8868 0.7939
kanzig2023unequal Carbon price 2000Q4 - 2019Q4 16.0719 0.188
CALDARA20191 Oil demand 2000Q4 - 2015Q4 12.0753 0.4396
CALDARA20191 Oil supply 2000Q4 - 2015Q4 10.8998 0.5375
BH2019 Oil supply 2000Q4 - 2025Q2 9.2402 0.6823
BH2019 Oil consumption demand 2000Q4 - 2025Q2 11.9281 0.4515
BH2019 Oil inventories 2000Q4 - 2025Q2 20.1281 0.0647
Baumeister2023 Oil price expectation 2000Q4 - 2023Q1 6.1135 0.9102
kilian2009not Oil supply 2000Q4 - 2009Q4 5.7186 0.9296
kilian2008exogenous Oil Supply 2000Q4 - 2016Q4 8.2033 0.7690
@@
Energy Market Responses to a Gas Supply Shock
@@
Figure 6 shows the impulse response functions from the VAR estimated by using the external instrument approach. As in kanzig2021macroeconomic, we can interpret the impulse response functions as elasticities, as all the variables in the VAR are in logs. We normalize the shock to increase the gas price, measured by the TTF, by 1%. We can see that price variables immediately react to the shock. The TTF responds positively and the response reverts to 0 in a bit more than an year. The increase in the price of natural gas is then transmitted to the national electricity price, measured by the PUN, which increases strongly, peaking within a few months and then slowly reverting. This confirms pass-through from gas to electricity: gas-fired plants set the marginal price, so the electricity market mirrors gas dynamics with a short lag. About half of the increase in the gas price is transmitted to the electricity price on impact. After a few months, gas inventories decrease of about 0.1%. Higher gas prices coincide with withdrawals from storage before restocking once prices ease. Gas quantity initially do not respond to the shock. After 10 months, we observe a small drop in gas quantity followed by a gradual recovery. Industrial production is not affected by the shock, consistently with its transitory nature.
Gas supply shock
@@
[Uncaptioned image] Figure 6: Impulse response functions of energy market variables to a natural gas supply shock. The figure reports the impulse response functions of natural gas price, gas inventories, gas quantity, industrial production, PUN and world industrial production to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: Ministero dell’ambiente e della sicurezza energetica, Gestore Mercati Energetici, ISTAT, Refinitiv Workspace, BH2019. Sample: 2002:Q4–2025:Q2.
@@ 3.2 Natural gas supply shocks and firms’ behavior
In order to evaluate how natural gas supply innovations affect the pricing decisions of firms, we estimate model shown in Section 3.1 and we extract the identified structural shock. Then, we estimate a proxy-internalizing VAR, placing the shock of interest as the first variable in the system and imposing recursivity (BRULU2022). [^8] According to PMW2021, the advantage of this approach over the equivalent SVAR-IV method of SW2018, besides its simplicity, is that the estimated impulse responses are valid even if the shock of interest is non-invertible. Apart from our estimated innovation, the model contains our index of inflation uncertainty, the average firms’ annual rate of change in their own prices and the average expected change in their prices over the next year, the annual rate of change in their assessment about order-book levels and about the stock of finished products and their inflation expectations. Firms expectations from SIGE are measured at quarterly frequency, at the final month of each quarter. In order to preserve the proper time alignment, we follow AB2023 and we take the end of the quarter values of the gas supply shock. Our sample spans from 2005Q4 until 2025Q2.
@@
[Uncaptioned image] Figure 7: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of inflation uncertainty index, firms’ expected and realized annual price change, firms’ assessed order-book levels and stock of finished products and annual inflation expectations to a natural gas supply shock. Shaded areas represent 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT, European Commission - ECFIN. Sample: 2005:Q4–2025:Q2.
Due to the limitations of our sample size, we resort to Bayesian methods for recovering the impulse response functions. We adopt the hierarchical approach of giannone2015prior for the optimization of the hyperparameters and the lenza2022estimate correction, in order to account for the huge volatility of shocks during Covid. As all our variables are stationary, we estimate the model by shrinking the autoregressive term to 0. We set the number of lags to 4. Figure 7 reports the generalized impulse response functions to a shock that increases the TTF by 1%. All dependent variables in this specification are expressed in annual percentage changes. The results are consistent with a cost-push disturbance. On impact, the annual rate of change in firms’ prices increases by 5% and the effect fades in a bit more than one year. Expectations about future prices display a similar but shorter-lived pattern. The year on year change jumps higher with respect to price realizations on impact, and reverts to 0 in half an year. Inflation uncertainty increases sharply for about one year, suggesting that the shock not only pushes up the level of prices but also complicates firms’ forecasting. Inflation expectations move in the same direction and with a similar scale as price expectations, but the increase in their annual growth rate is as persistent as the one for firms’ realized prices.
@@
adolfsen2024gas argue that the pass-trough of a natural gas supply shock in the Euro Area is characterized by a non linearity in the level of tightness in the labor market, with effects that are considerably magnified when the economy operates near full capacity. We show that the pass-through of natural gas supply shocks is also characterized by substantial non-linearities related firms’ ability to accurately forecast inflation. When inflation uncertainty increases, it becomes more difficult for firms to form an expectation about the future level of prices. This affects the way in which they respond to the shock of the price of natural gas. Under high uncertainty scenarios, the pass through is larger and firms persistently increase their prices. When instead uncertainty about future inflation is low they reduce their prices in order to preserve their market shares, anticipating the recessionary effects of the shock. We illustrate this by adopting state dependent local projections (ramey2018government). We determine states of high and low firms’ inflation uncertainty using the index constructed using the firm level micro data from the SIGE survey by adopting the approach of binder2017measuring. Following F2021 we scale the uncertainty index for the average inflation expectations of firms, in order to account for high inflationary periods, and we smooth it by applying a 4 period backward looking weighted moving average filter. We assume that the transition between states of high and low probabilities of inflation uncertainty is governed a logistic function $Z(\hat{\Delta}_{t-1})\in[0,1]$ of the lagged index, defined as:
@@
Where $Z(\hat{\Delta}_{t-1})$ indicate the probability of being in the state of high inflation uncertainty, $\bar{Z}\equiv 1-Z(\hat{\Delta}_{t-1})$, $\textit{controls}(x,\bar{S})$ is a set of control variables including the interaction between $w^{gas}_{t}$ and the probability of being in the tight labor market state $S(\hat{\Delta}_{t-1})$ and lags of the shock and of the dependent variable $x_{t}$. [^10] In our baseline specification, we include two lags of the dependent variable and two lags of the shock to preserve degrees of freedom. Results are robust to specifications with four lags of both variables. $D_{t}$ is a dummy variable that assumes value of 1 in correspondence of 2020Q2 and then decays exponentially over the subsequent quarters. The estimated impulse response function are reported in Figure 9.
Gas supply shock
Figure 9: Firms annual price change state-dependent response to a gas supply shock. The figure shows the state dependent response of firms’ prices to a gas supply shock. Estimation follows the lag-augmentation method of MOPM2021. The number of lags of the dependent variable, the control variables and the shock in each state are determined by minimizing the AIC. Shaded bands represent 68% confidence intervals based on Newey–West standard errors. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 2007Q1–2025Q2.

### v4

@@ 2 October 2022 Russia halts gas deliveries to Italy trough the TAG pipeline
17 November 2022 Russia permanently suspends gas flows via Austria
@@
We end up isolating 17 surprises which are reported in Table 1. We construct the daily instrument as the log-difference of the Dutch TTF day-ahead future between the day of the event and the last trading day preceding it. Following the argument in Kanzig2020, if the risk premium remains constant within this narrow daily window, the resulting price movement can be interpreted as a news-induced revision in the expected price of natural gas. Then, we aggregate the daily instrument at the monthly level by taking the sum within the month.
@@
We specify a monthly VAR including six variables. We take the day ahead Dutch TTF as a measure of the spot price of natural gas. We make this price real by deflating it by a measure of the Italian HICP we built by excluding the price of natural gas. Then, in the spirit of adolfsen2024gas, we include the available gas quantities, gas inventories and industrial production for Italy. We also control for the world industrial production and for the real ”Prezzo Unico Nazionale” (PUN) as a measure of the Italian spot electricity price. [^8] Data about gas quantities and gas inventories are provided by the Italian Ministry of Environment and Energy Security. Gas quantities are defined as national production, plus imports minus exports. Gas inventories are given by the cumulative sum of the variation in inventories, given a base value of 7000 millions of standard cubic meters. Both gas quantities and gas inventories are seasonally adjusted by using the X-13 ARIMA seasonal adjustment method. Data about the PUN are provided by Gestore Mercati Energetici (GME). World industrial production is provided by BH2019. Monitoring the response of the PUN to a gas supply shock is informative because, in a uniform-price electricity market with gas-fired units often setting the marginal cost, like the Italian one, changes in the TTF day-ahead price feed into generators’ bids and hence the Italian spot price with short lags; the PUN therefore reflects downstream pass-through from gas to electricity price. Controlling for world industrial production purges the gas-price equation of global demand pressure, sharpening identification of a gas supply shock. The VAR is estimated in log levels with a constant from 2004M1 until 2025M6. The lag order is set at 12. We assess instrument relevance by regressing the VAR residuals of the TTF price on the instrument and computing the corresponding F-statistic, following olea2021inference. An F-value above 10 indicates a strong instrument. Despite the limited number of events in our narrative surprise series, the instrument is strong: in our case the F-statistic is 22.08.
Figure 8: Gas supply shock. The figure shows the natural gas supply shock obtained from the IV-VAR (left) and the corresponding correlogram (right). Sample: 2005:M4–2025:M6.
We also conduct the F-test allowing for heteroskedasticity, which delivers an F-statistic of 17.66.999The instrument explains around 8% of the TTF residual. Figure 8 reports the structural shock extracted from the IV-VAR and the corresponding correlogram.
@@
We test the exogeneity of our shock with respect to oil and energy-related shocks identified in the literature using a Wald test of joint significance. Specifically, we regress our shock measure on the contemporaneous and 12 lagged values of each structural innovation considered, as well as on 12 of its own lags. We consider the oil price news shock of Kanzig2020, the carbon price shock of kanzig2023unequal, the oil demand shocks of caldara2019oil and BH2019, the oil supply,inventories and demand shock of BH2019, the oil price expectation shock of Baumeister2023, and the oil supply shocks of caldara2019oil, kilian2009not and kilian2008exogenous. Results are reported in Table 2. The Wald tests do not reject the null hypotesis of no contamination for any proxy, supporting exogeneity of our shock.
@@
Kanzig2020 Oil price news 2005M4 - 2024M6 16.9221 0.1525
kanzig2023unequal Carbon price 2005M4 - 2019M12 13.2596 0.3505
caldara2019oil Oil demand 2005M4 - 2015M11 10.4933 0.5728
caldara2019oil Oil supply 2005M4 - 2015M11 14.1466 0.2914
BH2019 Oil supply 2005M4 - 2025M5 9.2402 0.6823
BH2019 Oil consumption demand 2005M4 - 2025M5 12.3471 0.4182
BH2019 Oil inventories 2005M4 - 2025M5 20.1281 0.0647
Baumeister2023 Oil price expectation 2005M4 - 2023M3 7.8753 0.7947
kilian2009not Oil supply 2005M4 - 2009M12 4.1983 0.9796
kilian2008exogenous Oil Supply 2005M4 - 2016M12 11.2073 0.5112
@@
Gas supply shock transmission
@@
Figure 9 shows the impulse response functions from the VAR estimated using the external instrument approach. As in kanzig2021macroeconomic, we can interpret the impulse response functions as elasticities, since the model is estimated in log levels. We normalize the shock to increase the gas price, measured by the TTF, by 1%. We observe that price variables immediately react to the shock. The TTF responds positively, and the response reverts to 0 in about one year. The increase in the price of natural gas is then transmitted to the national electricity price, measured by the PUN, which increases strongly, peaking within a few months and then slowly reverting. This confirms the pass-through from gas to electricity: gas-fired plants set the marginal price, so the electricity market mirrors gas dynamics with a short lag. About half of the increase in the gas price is transmitted to the electricity price in the second month after the shock. After a few months, gas inventories decrease. Higher gas prices coincide with withdrawals from storage before restocking once prices ease. Gas quantities initially do not respond to the shock on impact. After 20 months, we observe a small drop in gas quantities followed by a gradual recovery. Industrial production is not affected by the shock, consistent with its transitory nature. To verify the robustness of our results, we estimate the model including the measure of geopolitical risk by caldara2022measuring. [^10] Adding this additional variable to the model does not affect the strength of our instrument. The F statistic is 18.68 and the heteroskedasticity robust F statistic is instead 17.58. Impulse response functions for the model that includes the measure of geopolitical risk are reported in appendix B. Our baseline results are not affected by the inclusion in the model of this additional variable. This evidence suggests that our instrument is not merely capturing a generic geopolitical risk shock, but predominantly identifies shocks to the physical supply of natural gas relevant for Italy.
@@
Figure 9: Impulse response functions of gas market variables, electricity prices and industrial production to a natural gas supply shock. The figure reports the impulse response functions of natural gas price, gas inventories, gas quantity, industrial production, PUN and world industrial production to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: Ministero dell’ambiente e della sicurezza energetica, Gestore Mercati Energetici, ISTAT, Refinitiv Workspace, BH2019. Sample: 2004:M1–2025:M6.
To obtain a deeper understanding of how the shock propagates through the economy, we follow kanzig2021macroeconomic and re-estimate the external-proxy VAR by replacing the electricity price with a set of alternative variables, one at a time. We begin with producer price indices, distinguishing among capital, energy, and intermediate goods. Figure 10 presents the results.
Energy producer prices display a clear and statistically significant hump-shaped response: the shock raises energy input costs by roughly 0.2 – 0.4 percent over the first ten months, after which the effect gradually dissipates. Compared with the spot electricity price, the producer price index for energy exhibits a smoother and more persistent hump-shaped adjustment, consistent with the slower pass-through of energy costs through regulated tariffs, contracts and distribution margins. The response of capital-goods producer prices is modest and highly persistent, but only marginally significant. Finally, the producer price index for intermediate goods shows no statistically meaningful reaction to the shock.
Refer to captionRefer to captionRefer to caption
Figure 10: Impulse response functions of producer price indexes to a natural gas supply shock. The figure reports the impulse response functions of energy, capital goods and intermediate goods producer price indexes to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: ISTAT.
Then, we consider a set of fuel prices provided by the Italian Ministry of Environment and Energy Security. [^11] The prices are considered net of excise and vat, and have been deflated using the italian HICP. The full set of fuel prices is available at https://sisen.mase.gov.it/dgsaie/prezzi-mensili-carburanti. All the responses for fuel prices appear to be positive, but most of them are barely significant. The only fuel that shows a statistically significant reaction is LPG. Its impulse response features a clear and persistent increase. This result is consistent with LPG being partially linked to natural gas markets and therefore more directly exposed to gas-related supply disruptions. However, this fuel has a limited importance for firms. For industrial fuels, neither light nor dense fuel oil exhibits a statistically significant reaction. Light fuel oil shows a positive median hump, but confidence bands remain wide and always include zero. Overall, the evidence indicates that LPG is the only fuel price responding significantly to gas supply shocks, while all other fuels—whether retail or industrial—show no statistically robust adjustment. This confirms that the transmission of gas supply shocks through fuel markets is generally limited in Italy and does not represent a major channel of cost pass-through for firms.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 11: Impulse response functions of real fuel prices to a natural gas supply shock. The figure reports the impulse response functions of the real price of gasoline, diesel fuel, heating diesel, LPG, dense and light fuel oil to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: Ministry of Environment and Energy Security.
To assess whether natural gas supply disruptions propagate through financial channels, we examine equity prices, sovereign yields, short-term interest rates, and a composite indicator of systemic stress (CISS). Results are reported in Figure 12. Equity prices fall on impact and remain below baseline for several months. The 10-year government bond yield moves up in the first part of the horizon, while short-term interest rates show a small and delayed increase. The CISS reacts only marginally and quickly returns to baseline. Overall, the only responses that are statistically significant at the 68% confidence level are the initial drop in equity prices and, to a lesser extent, the short-run increase in sovereign yields, while the movements in short-term rates and systemic stress remain mostly insignificant. Overall, the evidence points to a narrow transmission of natural gas supply shocks through the energy cost channel, but its spillovers to financial markets are limited.
Refer to captionRefer to captionRefer to caption
Figure 12: Impulse response functions of share prices, 10-year goverment bond yield, short term interest rate and CISS to a natural gas supply shock. The figure reports the impulse response functions of share prices, 10-year goverment bond yield and interest rate to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals, computed using a moving block bootstrap procedure. Source: OECD, EUROSTAT, ECB.
@@ 3.2 Natural gas supply shocks and firms’ behavior
In order to evaluate how natural gas supply innovations affect the pricing decisions of firms, we estimate model shown in Section 3.1 and we extract the identified structural shock. Then, we estimate a proxy-internalizing VAR, placing the shock of interest as the first variable in the system and imposing recursivity (BRULU2022). [^12] According to PMW2021, the advantage of this approach over the equivalent SVAR-IV method of SW2018, besides its simplicity, is that the estimated impulse responses are valid even if the shock of interest is non-invertible. Apart from our estimated innovation, the model contains our index of inflation uncertainty, the average firms’ annual rate of change in their own prices and the average expected change in their prices over the next year, their inflation expectations and our proxies for the intensive and extensive margins. Firms expectations from SIGE are measured at quarterly frequency, at the final month of each quarter, hence the estimated gas price shock is aggregated at a quarterly level by taking the sum within the quarter. The model is estimated from 2005Q1 until 2025Q2.
@@
[Uncaptioned image] Figure 13: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of the inflation uncertainty index, firms’ expected and realized annual price change, inflation expectations, intensive and extensive margins and industrial confidence index to a natural gas supply shock. Shaded areas represent 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 2005:Q1–2025:Q2.
Due to the limitations of our sample size, we resort to Bayesian methods to recover the impulse response functions. We adopt the hierarchical approach of giannone2015prior for the optimization of the hyperparameters. [^13] Since our variables in our BVAR do not exhibit any outlier in correspondence of the first and second quarter of 2020, we are not able to estimate the model by adopting the lenza2022estimate correction for accounting for increased Covid variability. Instead, we show that our IRFs are essentially unaffected when we estimate the model by completely removing the observations of 2020 from our sample. The lag-decay parameter of the Minnesota prior is treated as another hyperparameter and is estimated as well. As all our variables are stationary, we estimate the model by setting the prior mean on the coefficient of the first own lag and the prior mean of the sum-of-coefficients prior to 0. We set the number of lags to 4. Figure 13 reports the impulse response functions to a shock that increases the TTF by 1%. The results are consistent with a cost-push disturbance. On impact, inflation uncertainty increases sharply for about one year, suggesting that the shock not only pushes up the level of prices but also complicates firms’ forecasting. Firms’ prices exhibit a hump-shaped response, increasing persistently. Their response reverts to 0 in a little more than one year. The pass-through of input cost shocks to firms’ selling prices is however quantitatively limited. A gas supply shock that raises the real natural gas price by 1 percent translates into a peak increase of about 0.1 percentage points in firms’ annual price inflation. This is consistent with riggi2022price. One-year-ahead expectations about future prices display a shorter-lived increase: on impact, they increase by about 0.5 percentage points and revert to 0 within half a year. Firms’ one-year-ahead inflation expectations increase immediately on impact, showing a hump-shaped response. The response of the price margins indicates that the shock not only increases the proportion of firms that update their price; it also induces the price-adjusting firms to adjust more aggressively when they do update, as the intensive and extensive margins rise gradually. As a robustness, we compare the responses computed by adopting our baseline VAR with the ones obtained by adopting local projections. Results are reported in Figure 14. The impulse response functions delivered by the two models appear to be similar.
Gas supply shock
[Uncaptioned image] Figure 14: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of the inflation uncertainty index, firms’ expected and realized annual price change, inflation expectations, intensive and extensive margins and industrial confidence index to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals. Blue dotted lines represent the IRF computed by using our baseline VAR and the corresponding 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 2005:Q1–2025:Q2.
Our results are robust to using the Jurado2015 measure of firms’ inflation uncertainty we estimate by using the high dimensional dataset for the Italian Economy provided by barigozzi2024large. Also in this case, the response of firms’ inflation uncertainty appear to be similar both in qualitative and quantitative terms. The Responses of this alternative model specification are reported in appendix C of the Appendix.
Sectoral analysis
riggi2022price document the fact that the firm size and geographical area do not appear to be factors affecting the size pass-trough of a shock affecting input prices. Instead, they find that the firm sector appear to be a relevant factor. In order to quantify the differences in the pass trough of a gas supply shock across sector, we repeat the estimation of the model shown in the previous paragraph by using the expected prices and inflation, the price changes over the last year and the extensive and the intensive margins for the manufacturing and the services sectors separately. [^14] As no significant differences emerge in terms of uncertainty among these two sector, we use the aggregate index of uncertainty for the total of the firms. This allows us to analyze differences in the price setting behavior of firms belonging to these two sectors conditionally to a natural gas supply shock. Figure 15 reports the estimated impulse response functions of the price setting variables for the two sectors. First, we see that for manufacturing firms the shock causes a markedly stronger response across all price-setting dimensions. Both realized and expected prices rise more sharply and reach their peak sooner compared with services firms, indicating that manufacturing firms adjust more aggressively to higher input costs. The intensive margin shows a similarly pronounced reaction, with manufacturing firms increasing the size of their price changes more strongly in the short run. The extensive margin also displays significant expansion, suggesting that a larger share of manufacturing firms revise their prices following the shock.
For services firms, the profile is different. The response of their own prices is barely significant and lower in magnitude with respect to firms from manufacturing sector and their inflation expectations respond with lower magnitude on impact before growing sluggisly and aligning with the one from manifacturing sector. In terms of price setting margins, the repsonse of the intensive margin is not significant, while the number of service sector firms updating their prices display a positive hump shaped response.
Overall, the results indicate that manufacturing firms react more promptly and forcefully to natural gas supply shocks, especially terms of the size of the adjustments, while the likelihood of adjusting prices is similar across the two sectors. These differences emerge despite the fact that the impulse response functions implied by those two alternative models do not imply any difference in terms of the response of inflation expectations to the shock. This finding is coherent with riggi2022price which find that inflation expectations are not an important determinant of firms’ price setting behavior, which instead are more influenced by the firms’ expected path of their own prices.
Figure 15: Impulse response functions of firms’ related variables to a natural gas supply shock. The figure reports the impulse response functions of the inflation uncertainty index, firms’ expected and realized annual price change, inflation expectations, intensive and extensive margins and industrial confidence index to a natural gas supply shock. Shaded areas represent 68% and 90% confidence intervals. Blue dotted lines represent the IRF computed by using our baseline VAR and the corresponding 68% confidence intervals. Source: Bank of Italy, Survey on inflation and growth expectations. Sample: 2005:Q1–2025:Q2.
@@
adolfsen2024gas argue that the pass-trough of a natural gas supply shock in the Euro Area is characterized by a non linearity in the level of tightness in the labor market, with effects that are considerably magnified when the economy operates near full capacity. We show that the pass-through of natural gas supply shocks is also characterized by substantial non-linearities related firms’ ability to accurately forecast inflation. When inflation uncertainty increases, it becomes more difficult for firms to form an expectation about the future level of prices. This affects the way in which they respond to the shock. Under high uncertainty scenarios, the pass through is larger and firms persistently increase their prices. When instead uncertainty about future inflation is low they reduce their prices in order to preserve their market shares, anticipating the recessionary effects of the shock. We illustrate this by adopting state dependent local projections (ramey2018government). We determine states of high and low firms’ inflation uncertainty using the index constructed using the firm level micro data from the SIGE survey by adopting the approach of binder2017measuring. Following F2021 we scale the uncertainty index for the average inflation expectations of firms, in order to account for high inflationary periods, and we smooth it by applying a 4 period backward looking weighted moving average filter. We assume that the transition between states of high and low probabilities of inflation uncertainty is governed a logistic function $Z(\hat{\Delta}_{t-1})\in[0,1]$ of the lagged index, defined as:
@@
Where $Z(\hat{\Delta}_{t-1})$ indicate the probability of being in the state of high inflation uncertainty, $\bar{Z}\equiv 1-Z(\hat{\Delta}_{t-1})$, $\textit{controls}(x,\bar{S})$ is a set of control variables including the interaction between $w^{gas}_{t}$ and the probability of being in the tight labor market state $S(\hat{\Delta}_{t-1})$ and lags of the shock and of the dependent variable $x_{t}$. [^16] In order to preserve degrees of freedom, we set the number of lags by minimizing the AIC at horizon $h=0$, allowing for a maximum of four lags. $D_{t}$ is a dummy variable that assumes value of 1 in correspondence of 2020Q2 and then decays exponentially over the subsequent quarters. The estimated impulse response function are reported in Figure 17.
Figure 17: Firms annual price change state-dependent response to a gas supply shock. The figure shows the state dependent response of firms’ prices to a gas supply shock. Estimation follows the lag-augmentation method of MOPM2021. The number of lags of the dependent variable, the control variables and the shock in each state are determined by minimizing the AIC at horizon $h=0$. Shaded bands represent 68% and 90% confidence intervals based on Newey–West standard errors. Source: Bank of Italy, Survey on inflation and growth expectations, EUROSTAT. Sample: 2007Q1–2025Q2.